OpenStudy (dtan5457):

suppose p and q are positive #'s which log9(p)=log12(q)=log16(p+q) what is q/p?

myininaya (myininaya):

\[\log_9(p)=\log_{12}(q)=\log_{16}(p+q)\]?

OpenStudy (dtan5457):

yes

myininaya (myininaya):

so maybe we can do something with this: \[y=\log_9(p) \implies 9^y=p \\ y=\log_{12}(q) \implies 12^{y}=q \\ y=\log_{16}(p+q)=\log_{16}(9^y+12^y) \implies 16^y=9^y+12^y \\....\] just putting on some first thoughts

OpenStudy (dtan5457):

should i guess values for y now?

myininaya (myininaya):

y is hard to guess

OpenStudy (astrophysics):

Try change of base formula should be easier

OpenStudy (anonymous):

i seem to remember one like this \[\log_9(p)=\frac{\log(p)}{2\log(3)}\] for example

OpenStudy (astrophysics):

Yes, exactly \[\log_9 p = \frac{ \log p }{ \log 9 }\] \[\log _{12}q = \frac{ \log q }{ \log_{12} }\] and note \[\log_9p = \log_{12} q\] so you have a ratio and you can figure out q/p

OpenStudy (astrophysics):

\[\huge \log_a x= \frac{ \log_b x }{ \log_b a }\] change of base

OpenStudy (dtan5457):

i can see that but how does that help me solve for p or q?

myininaya (myininaya):

oh I think I have something playing with what I had above \[\frac{p}{q}=\frac{3^y}{4^y} \implies 4^y \frac{p}{q}=3^y \text{ or if squaring both sides } \\ \implies 4^{2y}(\frac{p}{q})^2=3^{2y} \\ \text{ so anyways I want to write } 16^y=9^y+12^y \text{ in terms of } 3^y \text{ and } 4^y \\ \text{ so we have } 4^{2y}=3^{2y}+3^y 4^y\] write the equation 4^y and p/q using the two equations I just mentioned

myininaya (myininaya):

you will end up with an equation just in terms of p/q which you could replace with u the equation I have is a quadratic

myininaya (myininaya):

And then once you find p/q you can just flip it to obtain q/p

myininaya (myininaya):

so in case you don't see you only have two things to replaced the 3^(2y) thing and the 3^y

myininaya (myininaya):

you are writing my equation above in terms of 4^y and p/q \[3^y=4^y \frac{p}{q} \\ 3^{2y}=4^{2y}(\frac{p}{q})^2\]

OpenStudy (dtan5457):

Where did you get 3y and 4y?

myininaya (myininaya):

you mean 3^y and 4^y?

myininaya (myininaya):

from my equations above: \[y=\log_9(p) \implies 9^y=p \\ y=\log_{12}(q) \implies 12^{y}=q \\ y=\log_{16}(p+q)=\log_{16}(9^y+12^y) \implies 16^y=9^y+12^y \\....\] \[\frac{p}{q}=\frac{9^y}{12^y}=\frac{3^{y} 3^{y}}{3^{y} 4^{y}}=\frac{3^y}{4^y}\]

OpenStudy (dtan5457):

im confused about what to do after we get to 4^2y=3^2y+3^y(4^y)

myininaya (myininaya):

you make these replacements yet: \[3^y=4^y \frac{p}{q} \\ 3^{2y}=4^{2y}(\frac{p}{q})^2\]

ganeshie8 (ganeshie8):

\[p+q = 16^x = \dfrac{(16*9)^x}{9^x} = \dfrac{12^x12^x}{9^x} = \dfrac{q*q}{p}\]

ganeshie8 (ganeshie8):

looks we can't avoid a quadratic...

myininaya (myininaya):

\[4^{2y}=4^{2y}(\frac{p}{q})^2+4^y \frac{p}{q}4^y \\ 4^{2y}=4^{2y}(\frac{p}{q})^2+4^{2y}\frac{p}{q} \\ 1=(\frac{p}{q})^2+\frac{p}{q}\]

OpenStudy (dtan5457):

ok i replaced it

myininaya (myininaya):

@ganeshie8 way is faster getting there

OpenStudy (dtan5457):

@ganeshie8 can you breifly explain what you did

OpenStudy (dtan5457):

you put log16(p+q) into exponential form and im not sure what you did after that

myininaya (myininaya):

he multiply by a pretty one called 9^x/9^x

myininaya (myininaya):

then rearrange some things to see p and q where p is 9^x and q is 12^x

ganeshie8 (ganeshie8):

@dtan5457 The equation \(\log_b a = \color{red}{x}\) is same as the equaiton \(a=b^{\color{red}{x}}\)

ganeshie8 (ganeshie8):

\(a=b^{\color{red}{x}}\) Notice here, \(\color{red}{x}\) is the exponent. \(\color{red}{\text{logarithm}} = \color{red}{\text{exponent}}\)

ganeshie8 (ganeshie8):

\[\log_{16}(p+q) =x\] is same as the equation \[ p+q=16^x\]

OpenStudy (dtan5457):

yes i get that

ganeshie8 (ganeshie8):

\[\log_9(p)=\log_{12}(q)=\log_{16}(p+q)=x\] From above equation, do we have : \(p = 9^x\) \(q=12^x\) \(p+q=16^x\) ?

OpenStudy (dtan5457):

yes @myininaya had showed that earlier

ganeshie8 (ganeshie8):

Now take a look at my first reply

ganeshie8 (ganeshie8):

This one : \[p+q = 16^x = \dfrac{(16*9)^x}{9^x} = \dfrac{12^x12^x}{9^x} = \dfrac{q*q}{p}\]

OpenStudy (dtan5457):

i get the last 2 fractions but the first one im not sure what you did

OpenStudy (dtan5457):

why 16 x 9?

ganeshie8 (ganeshie8):

Ahh that is hard to explain. We have introduced a new variable \(x\) in order to eliminate the nasty logs. We are trying to eliminate \(x\) too and get an equation in terms of \(p,q\) only.

ganeshie8 (ganeshie8):

In pursuit of that, we're trying to express 16 in terms of 9 and 12

ganeshie8 (ganeshie8):

Notice \(16=4*4 = \dfrac{4*4*3*3}{3*3} = \dfrac{12*12}{9}\)

OpenStudy (dtan5457):

theres no way i would of thought of that but yeah now i see what you did

OpenStudy (dtan5457):

now that we know q^2/p is equal to 16^x what do we do

ganeshie8 (ganeshie8):

That is not just the way to do it. There are multiple ways to do this. You just use whatever the idea that occurs to you

ganeshie8 (ganeshie8):

\[p+q = 16^x = \dfrac{(16*9)^x}{9^x} = \dfrac{12^x12^x}{9^x} = \dfrac{q*q}{p} \] So we have \[p+q = \dfrac{q^2}{p}\]

ganeshie8 (ganeshie8):

logs and x are gone

ganeshie8 (ganeshie8):

divide \(q\) through out and let \(\frac{q}{p}=t\)

OpenStudy (dtan5457):

(p/q)+1=q/p?

OpenStudy (dtan5457):

or t+1=q/p

ganeshie8 (ganeshie8):

(p/q)+1 = q/p lettinh q/p = t, we get 1/t + 1 = t

ganeshie8 (ganeshie8):

which is same as 1 + t = t^2 you can solve it

OpenStudy (dtan5457):

(1+sqrt5)/2?

OpenStudy (dtan5457):

the minus is rejected yes?